# A Survey on RAG Meets LLMs: Towards Retrieval-Augmented Large Language Models

### v1

@@ Abstract.
As one of the most advanced techniques in AI, Retrieval-Augmented Generation (RAG) techniques can offer reliable and up-to-date external knowledge, providing huge convenience for numerous tasks. Particularly in the era of AI-generated content (AIGC), the powerful capacity of retrieval in RAG in providing additional knowledge enables retrieval-augmented generation to assist existing generative AI in producing high-quality outputs. Recently, large Language Models (LLMs) have demonstrated revolutionary abilities in language understanding and generation, while still facing inherent limitations, such as hallucinations and out-of-date internal knowledge. Given the powerful abilities of RAG in providing the latest and helpful auxiliary information, retrieval-augmented large language models have emerged to harness external and authoritative knowledge bases, rather than solely relying on the model’s internal knowledge, to augment the generation quality of LLMs. In this survey, we comprehensively review existing research studies in retrieval-augmented large language models (RA-LLMs), covering three primary technical perspectives: architectures, training strategies, and applications. As the preliminary knowledge, we briefly introduce the foundations and recent advances of LLMs. Then, to illustrate the practical significance of RAG for LLMs, we categorize mainstream relevant work by application areas, detailing specifically the challenges of each and the corresponding capabilities of RA-LLMs. Finally, to deliver deeper insights, we discuss current limitations and several promising directions for future research.
Retrieval-Augmented Generation (RAG), Large Language Models (LLMs), Pre-training, Fine-tuning, In-context Learning, Prompting.
@@ 1. Introduction
As one of the most fundamental data mining techniques, retrieval aims to understand the input query and extract relevant information from external data sources (Kobayashi and Takeda, 2000; Singhal et al., 2001). It has found extensive application in various fields (Buttcher et al., 2016; Yin et al., 2016; O’Hare et al., 2016), such as search, question answering, and recommender systems. For instance, search engines (e.g., Google, Bing, and Baidu) are the most successful applications of retrieval in the industry; they can filter and retrieve the most relevant web pages or documents that can match a user’s query (Croft et al., 2010; Yin et al., 2016), enabling users to find the desired information effectively. Meanwhile, retrieval models, through effective data maintenance in external databases, can provide faithful and timely external knowledge, thereby serving vital functions in various knowledge-intensive tasks. Due to their powerful capacities, retrieval techniques have been successfully incorporated into advanced generative models in the era of AI-Generated Content (AIGC) (Li et al., 2023a; Wu et al., 2024; Sheynin et al., 2023; Zhang et al., 2023a). Notably, the integration of retrieval models with language models has given rise to Retrieval-Augmented Generation (RAG) (Lewis et al., 2020c), which has emerged as one of the most representative techniques in the field of generative AI, aiming to enhance the generation quality of text content (Li et al., 2023a; Lewis et al., 2020c; Borgeaud et al., 2022).
@@
Figure 1. Retrieval-Augmented Generation (RAG) meets Large Language Models (LLMs). When the user’s query is out-of-scope, e.g., unseen content in training data or the need for the latest information for the answer, LLMs might show inferior generation performance. With the help of RAG, LLMs can leverage additional relevant information from external data sources to enhance text generation capability.
@@
To advance generation models and enhance the generated results, RAG incorporates information or knowledge from external data sources, which serves as supplementary for the input query or the generated output (Min et al., 2020; Khandelwal et al., 2020). Specifically, RAG first invokes the retriever to search and extract the relevant documents from external databases, which are then leveraged as the context to enhance the generation process (Izacard and Grave, 2021b). In practice, RAG techniques are feasible and efficient to apply in various generation tasks with simple adaptation of the retrieval component, requiring minimal or even no additional training (Ram et al., 2023). Recent studies have demonstrated the great potential of RAG not only for knowledge-intensive tasks such as the Open-domain Question Answering (OpenQA) (Borgeaud et al., 2022; Guu et al., 2020; Petroni et al., 2020), but also for general language tasks (Khandelwal et al., 2020; He et al., 2021a; Xu et al., 2020), and various downstream applications (Wu et al., 2024; Liu et al., 2023).
@@
Recent years have witnessed the rapid development of pre-trained foundation models, particularly Large Language Models (LLMs), which have demonstrated impressive performance across various tasks (Chowdhery et al., 2023; Achiam et al., 2023), including recommender systems (Zhao et al., 2024a), molecule discovery (Li et al., 2023a), and report generation (Ding et al., 2024). The great success of LLMs can be technically attributed to the advanced architectures with billion-level parameters pre-training on a huge amount of training corpus from various sources. These technical improvements have given rise to the remarkable emergence capabilities of LLMs (Zhao et al., 2023b, 2024a), particularly in language understanding and generation, in-context learning, and others. For instance, GPT-FAR introduces detailed prompts to teach GPT-4 to perform image tagging, statistical analysis, and text analysis for multi-modal fashion report generation (Ding et al., 2024). LLMs also achieve promising performance in recommender systems by understanding users’ preferences towards items (Wang et al., 2024a; Zhao et al., 2024a). Despite the success, LLMs still suffer from intrinsic limitations (Zhao et al., 2024a, 2023b), such as the lack of domain-specific knowledge, the problem of “hallucination”, and the substantial computational resources for updating the models. These problems are particularly notable in domain-specific fields like medicine and law. For instance, a recent study has demonstrated that legal hallucinations are pervasive and disturbing, with hallucination rates ranging from 69% to 88% in response to specific legal queries for state-of-the-art LLMs (Dahl et al., 2024). Moreover, the challenges of tackling the hallucination problem become even harder due to the substantial computational resources required for fine-tuning LLMs with domain-specific or the latest data. This, in turn, significantly hinders the widespread adoption of LLMs in various real-world applications.
@@
To address these limitations, recent efforts have been made to take advantage of RAG to enhance the capabilities of LLMs in various tasks (Shi et al., 2023; Khandelwal et al., 2020; Borgeaud et al., 2022; Izacard and Grave, 2021a), especially those demanding high for the latest and reliable knowledge such as Question Answer (QA), AI4Science, and software engineering. For example,  Lozano et al. (2023) introduces a scientific-specific QA system based on retrieving scientific literature dynamically. MolReGPT leverages RAG to enhance the in-context learning ability of ChatGPT for molecular discovery (Li et al., 2023a). As illustrated in Figure 1, an LLM-based dialog system will not be able to answer well for out-of-scope queries. In comparison, with the help of RAG to retrieve relevant knowledge from external data sources and integrate it into the process of generation, the dialog system succeeds in giving correct answers to the user. Given the remarkable progress in advancing LLMs with RAG, there is an imperative need for a systematic review of recent advances in Retrieval-Augmented Large Language Models (RA-LLM).
@@
This survey aims to provide a comprehensive overview of the retrieval-augmented large language models, i.e., RA-LLMs, by summarizing representative methods from the aspects of RA-LLMs’ architecture, training, and applications. More specifically, following a brief introduction to the background knowledge of LLMs in Section 2, we review existing research from several primary perspectives of RA-LLMs in terms of retrieval, generation, and augmentation in Section 3, as well as the necessity and application frequency of retrieval in RAG. Then, we summarize the main training techniques of RA-LLMs in Section  4 and various RA-LLMs applications in Section  5. Finally, in Section 6, we discuss key challenges and potential directions for future exploration.
@@ 2.1. Large Language Models (LLMs)
Recently, the significant breakthrough of LLMs has revolutionized the field of artificial intelligence (Zhao et al., 2023b; Brown et al., 2020; Fan et al., 2024). The advanced LLMs are typically pre-trained on extensive data with billion-level parameters and have demonstrated the ability to understand and generate human-like text, leading to advancements in various natural language processing tasks such as text generation and information retrieval (Zhao et al., 2023b, 2024a). LLMs can be adapted to a variety of downstream tasks by fine-tuning them on specific datasets, allowing them to specialize in particular domains or applications. In general, most existing LLMs can be broadly divided into three main categories: Encoder-only, Decoder-only, and Encoder-Decoder models.
@@ 2.2.2. In-Context Learning (ICL)
Despite its effectiveness, ICL usually relies heavily on the quality of the provided demonstrations, which may lead to the generation of sub-optimal outputs. Even worse, ICL may not have enough necessary information or prior knowledge to guide the LLMs in generating accurate responses. To address the aforementioned limitations of ICL, more recent studies introduce Retrieval-Augmented Generation (RAG) technologies (Lewis et al., 2020c; Ram et al., 2023; Shi et al., 2023). By integrating retrieval with generation, RAG models provide a promising direction for enhancing the performance and adaptability of LLMs across various tasks.
@@ 3. Retrieval-Augmented Large Language Models (RA-LLMs)
The RAG framework in the era of LLMs generally consists of three major processes of retrieval, generation, and augmentation, as well as the mechanism to determine whether the retrieval is needed. In this section, we will introduce important techniques involved in each component.
Figure 2. Illustration of the basic Retrieval-Augmented Large Language Models (RA-LLMs) framework for a specific QA task, which consists of three main components: retrieval, augmentation, and generation. Retrieval may have different procedures with various designs, which optionally includes pre-retrieval and post-retrieval processes. The retrieved documents are further leveraged in generation with the augmentation module, which may have different designs based on the integration stage in the generation model.
@@ 3.1.1. Retriever Type
Retrieval methods can be generally categorized into two types: sparse and dense, based on the information encoding methods. Sparse retrieval is word-based and applied in text retrieval mostly, while dense retrieval embeds queries and external knowledge into vector spaces and is easily applied to various data formats.
@@
As a straightforward approach, sparse retrieval, e.g., TF-IDF and BM25 (Sparck Jones, 1972; Robertson et al., 2009), usually relies on inverted index matching along with the raw data input. For example, many studies directly apply BM25 for passage-level retrieval to facilitate their RAG (Chen et al., 2017; Ram et al., 2023; Zhong et al., 2022; Jiang et al., 2023; Zhou et al., 2022; Xu et al., 2023b), where passages are specifically represented as a bag of words and ranked based on term and inverse document frequencies (Izacard and Grave, 2021b). On top of offering supplementary to enhance the input of the generator, sparse retrieval has also been used to find examples as demonstrations for ICL for RA-LLMs (Ye et al., 2023b; Luo et al., 2023b; Rubin et al., 2022; Agrawal et al., 2023; Sia and Duh, 2023). The main limitation of applying sparse retrieval in RAG is its no-training nature, which makes the retrieval performance heavily rely on the quality of database construction and query generation. Moreover, such fixed term-based methods only support similarity retrieval, while cannot be adapted for other retrieval considerations demanding in LLM applications, such as the diversity (Drozdov et al., 2022).
@@
Dense retrieval, on the contrary, embeds the query and documents into continuous vector space with certain criteria, for example, semantic similarity (Karpukhin et al., 2020). Dense retrieval methods can often be trained, therefore hold more flexibility and potential in adaptation. As the key component of dense retriever, the embedding models have delicately different designs in existing RAG models. A simple design (Yogatama et al., 2021; Khandelwal et al., 2020; Lewis et al., 2020a; Wu et al., 2022) is to directly use a part of the generation model as the embedding layer of the retriever, which might be able to enhance the alignment between the retrieval and generation processes. BERT-based backbone (Devlin et al., 2019) is widely applied in retrieval models. One common retriever design is to construct two-stream encoders with BERT structure (one encoder for the query and the other for the documents), which is also called bi-encoder (Wu et al., 2020; Shi et al., 2023). Early-stage RAG methods tend to freeze (Borgeaud et al., 2022; Ram et al., 2023) or partially freeze (Lewis et al., 2020c) the parameters of the retriever to perform general-level relevant knowledge extraction and pay more attention to the knowledge leveraging and generator fine-tuning. Large-scale specialized pre-training further enhances RAG models to excel in more knowledge-intensive tasks. One typical success is Dense Passage Retriever (DPR) (Karpukhin et al., 2020), which uses a BERT-based backbone and is pre-trained specifically for the OpenQA task with question-answer pair data. DPR has shown strong capacity as a pre-trained retriever, facilitating many RAG models to succeed in various downstream tasks (Lewis et al., 2020c; Izacard and Grave, 2021b; Siriwardhana et al., 2023; Singh et al., 2021; Shi et al., 2023). It has also been regarded as the first step in the RAG paradigm for improving the performance of LLMs, which may further enhance the alignment of the embeddings between queries and relevant textual data through fine-tuning (Cheng et al., 2023). A recent study (Reichman and Heck, 2024) has also discovered that DPR training decentralizes how knowledge is stored in the network, creating multiple access pathways to the same information. With effective fine-tuning, bi-encoder retrievers are also applied widely in ICL-based RAG (Rubin et al., 2022; Poesia et al., 2022; Lu et al., 2023; Ye et al., 2023b; Milios et al., 2023; Li and Qiu, 2023). Specifically, they have been more often used for sentence embedding similarity-based retrieval, as well as for some special requirement in ICL, such as diverse example retrieval (Ye et al., 2023b).
@@
Another stream of widely applied dense retrievers in RA-LLMs has one-encoder structures, which may be based on Transformer, BERT or other off-the-shelf sequence modeling backbones. These one-encoder retrievers are generally pre-trained on large-scale unaligned documents by contrastive learning (Reichman and Heck, 2024), which may therefore excel for their versatility, meaning that they can transfer and generalize better to new domains or tasks. Such general-purpose pre-trained retrievers, e.g., Contriever (Gautier et al., 2022) and Spider (Ram et al., 2022), would be more flexible to use in LLMs targeting on various tasks and have demonstrated their effectiveness in many RA-LLM methods, such as In-Context RALM (Ram et al., 2023), Atlas (Izacard et al., 2023), Self-RAG (Asai et al., 2023b), and others (Shi et al., 2023). According to experimental results in existing studies (Yu et al., 2023a), for open-domain QA tasks, when cooperated with InstructGPT (Ouyang et al., 2022), applying general-purpose pre-trained retriever (Contriever) without fine-tuning achieves comparable performance to sparse retriever (BM25). However, they are both worse than the DPR model fine-tuned on target datasets, showing the effectiveness of fine-tuning on targeted tasks and data.

### v2

# A Survey on RAG Meeting LLMs: Towards Retrieval-Augmented Large Language Models

@@ Abstract.
As one of the most advanced techniques in AI, Retrieval-Augmented Generation (RAG) can offer reliable and up-to-date external knowledge, providing huge convenience for numerous tasks. Particularly in the era of AI-Generated Content (AIGC), the powerful capacity of retrieval in providing additional knowledge enables RAG to assist existing generative AI in producing high-quality outputs. Recently, Large Language Models (LLMs) have demonstrated revolutionary abilities in language understanding and generation, while still facing inherent limitations, such as hallucinations and out-of-date internal knowledge. Given the powerful abilities of RAG in providing the latest and helpful auxiliary information, Retrieval-Augmented Large Language Models (RA-LLMs) have emerged to harness external and authoritative knowledge bases, rather than solely relying on the model’s internal knowledge, to augment the generation quality of LLMs. In this survey, we comprehensively review existing research studies in RA-LLMs, covering three primary technical perspectives: architectures, training strategies, and applications. As the preliminary knowledge, we briefly introduce the foundations and recent advances of LLMs. Then, to illustrate the practical significance of RAG for LLMs, we systematically review mainstream relevant work by their architectures, training strategies, and application areas, detailing specifically the challenges of each and the corresponding capabilities of RA-LLMs. Finally, to deliver deeper insights, we discuss current limitations and several promising directions for future research. Updated information about this survey can be found at https://advanced-recommender-systems.github.io/RAG-Meets-LLMs/ [^1] This is the long version of the survey to appear at KDD 2024 (Fan et al., 2024a).
Retrieval-Augmented Generation (RAG), Large Language Model (LLM), Pre-training, Fine-tuning, In-context Learning, Prompting.
@@ 1. Introduction
As one of the most fundamental data mining techniques, retrieval aims to understand the input query and extract relevant information from external data sources (Kobayashi and Takeda, 2000; Singhal et al., 2001). It has found extensive application in various fields (Buttcher et al., 2016; Yin et al., 2016; O’Hare et al., 2016), such as search, question answering, and recommender systems. For instance, search engines (e.g., Google, Bing, and Baidu) are the most successful applications of retrieval in the industry; they can filter and retrieve the most relevant web pages or documents that can match a user’s query (Croft et al., 2010; Yin et al., 2016), enabling users to find the desired information effectively. Meanwhile, retrieval models, through effective data maintenance in external databases, can provide faithful and timely external knowledge, thereby serving vital functions in various knowledge-intensive tasks. Due to their powerful capacities, retrieval techniques have been successfully incorporated into advanced generative models in the era of AI-Generated Content (AIGC) (Li et al., 2023a; Wu et al., 2024; Sheynin et al., 2023; Zhang et al., 2023a). Notably, the integration of retrieval models with language models has given rise to Retrieval-Augmented Generation (RAG) (Lewis et al., 2020c), which has emerged as one of the most representative techniques in the field of generative AI, aiming to enhance the generation quality of text content with retrieved information (Li et al., 2023a; Lewis et al., 2020c; Borgeaud et al., 2022).
@@
Figure 1. Retrieval-Augmented Generation (RAG) meets Large Language Models (LLMs). When the user’s query is out-of-scope, e.g., unseen content in training data or the need for the latest information for the answer, LLMs might show inferior generation performance. With the help of RAG, LLMs can leverage additional relevant information from external database to enhance their text generation capability.
@@
To advance generation models and enhance the generated results, RAG incorporates information or knowledge from external data sources, which serves as supplementary for the input query or the generated output (Min et al., 2020; Khandelwal et al., 2020). Specifically, RAG first invokes the retriever to search and extract the relevant documents from external databases, which are then leveraged as the context to enhance the generation process (Izacard and Grave, 2021b). In practice, RAG techniques are feasible and efficient to apply in various generation tasks with simple adaptation of the retrieval component, requiring minimal or even no additional training (Ram et al., 2023). Recent studies have demonstrated the great potential of RAG not only for knowledge-intensive tasks such as the Open-domain Question Answering (OpenQA) (Borgeaud et al., 2022; Guu et al., 2020; Petroni et al., 2020; Shi et al., 2024), but also for general language tasks (Khandelwal et al., 2020; He et al., 2021a; Xu et al., 2020), and various downstream applications (Wu et al., 2024; Liu et al., 2023).
@@
Recent years have witnessed the rapid development of pre-trained foundation models, particularly Large Language Models (LLMs), which have demonstrated impressive performance across various tasks (Chowdhery et al., 2023; Achiam et al., 2023), including recommender systems (Zhao et al., 2024a), molecule discovery (Li et al., 2023a), and report generation (Ding et al., 2024). The great success of LLMs can be technically attributed to the advanced architectures with billion-level parameters pre-training on a huge amount of training corpus from various sources. These technical improvements have given rise to the remarkable emergence capabilities of LLMs (Zhao et al., 2023b, 2024a), particularly in language understanding and generation, in-context learning, and others. For instance, GPT-FAR introduces detailed prompts to teach GPT-4 to perform image tagging, statistical analysis, and text analysis for multi-modal fashion report generation (Ding et al., 2024). LLMs also achieve promising performance in recommender systems by understanding users’ preferences towards items (Wang et al., 2024a; Zhao et al., 2024a). Despite the success, LLMs still suffer from intrinsic limitations (Zhao et al., 2024a, 2023b), such as the lack of domain-specific knowledge, the problem of “hallucination”, and the substantial computational resources required for updating the models. These problems are particularly notable in domain-specific fields like medicine and law. For instance, a recent study has demonstrated that legal hallucinations are pervasive and disturbing, with hallucination rates ranging from 69% to 88% in responses to specific legal queries for state-of-the-art LLMs (Dahl et al., 2024). Moreover, the challenges of tackling the hallucination problem become even harder due to the substantial computational resources required for fine-tuning LLMs with domain-specific or the latest data. This, in turn, significantly hinders the widespread adoption of LLMs in various real-world applications.
@@
To address these limitations, recent efforts have been made to take advantage of RAG to enhance the capabilities of LLMs in various tasks (Shi et al., 2023; Khandelwal et al., 2020; Borgeaud et al., 2022; Izacard and Grave, 2021a), especially those demanding high for the latest and reliable knowledge such as Question Answer (QA), AI4Science, and software engineering. For example,  Lozano et al. (2023) introduces a scientific QA system based on retrieving scientific literature dynamically. MolReGPT leverages RAG to enhance the in-context learning ability of ChatGPT for molecular discovery (Li et al., 2023a). It is also been demonstrated that RAG can effectively reduce hallucinations in conversational tasks (Shuster et al., 2021; Xu et al., 2022). As illustrated in Figure 1, an LLM-based dialog system will not be able to answer well for out-of-scope queries. With the help of RAG to retrieve relevant knowledge from external database and integrate it into the process of generation, the dialog system succeeds in giving correct answers. Given the remarkable progress in advancing LLMs with RAG, there is an imperative need for a systematic review of recent advances in Retrieval-Augmented Large Language Models (RA-LLMs).
@@
This survey aims to provide a comprehensive overview of RA-LLMs by summarizing representative methods from the aspects of the architecture, training strategy, and application area respectively. More specifically, following a brief introduction to the background knowledge of LLMs in Section 2, we review existing research from several primary perspectives of RA-LLMs in terms of retrieval, generation, and augmentation in Section 3, as well as the necessity and application frequency of retrieval in RAG. Then, we summarize the main training techniques of RA-LLMs in Section  4 and various RA-LLMs applications in Section  5. Finally, in Section 6, we discuss key challenges and potential directions for future exploration.
@@ 2.1. Large Language Models (LLMs)
Recently, the significant breakthrough of LLMs has revolutionized the field of artificial intelligence (Zhao et al., 2023b; Brown et al., 2020; Fan et al., 2024b). The advanced LLMs are typically pre-trained on extensive data with billion-level parameters and have demonstrated the ability to understand and generate human-like text, leading to advancements in various natural language processing tasks such as text generation and information retrieval (Zhao et al., 2023b, 2024a). LLMs can be adapted to a variety of downstream tasks by fine-tuning them on specific datasets, allowing them to specialize in particular domains or applications. In general, most existing LLMs can be broadly divided into three main categories: Encoder-only, Decoder-only, and Encoder-Decoder models.
Figure 2. Representing RAG and RA-LLMs methods organized by their main design focus, proposed time and impact (shown by citation). Note that the year shown along with the model name matches with the reference while the arrangement of the these models is based on their first time being published.
@@ 2.2.2. In-Context Learning (ICL)
Despite its effectiveness, ICL usually relies heavily on the quality of the provided demonstrations (Liu et al., 2022b; Su et al., 2023), which may lead to the generation of sub-optimal outputs. Even worse, ICL may not have enough necessary information or prior knowledge to guide the LLMs in generating accurate responses. To address the aforementioned limitations of ICL, more recent studies introduce Retrieval-Augmented Generation (RAG) technologies (Lewis et al., 2020c; Ram et al., 2023; Shi et al., 2023). By integrating retrieval with generation, RAG models provide a promising direction for enhancing the performance and adaptability of LLMs across various tasks.
@@ 3. Retrieval-Augmented Large Language Models (RA-LLMs)
The RAG framework in the era of LLMs consists of several major processes: retrieval, generation, and augmentation, as well as the mechanism to determine whether the retrieval is needed. In this section, we will introduce important techniques involved in each process.
Figure 3. Illustration of the basic Retrieval-Augmented Large Language Models (RA-LLMs) framework for a specific QA task, which consists of three main components: retrieval, augmentation, and generation. Retrieval may have different procedures with various designs, which optionally includes pre-retrieval and post-retrieval processes. The retrieved documents are further leveraged in generation with the augmentation module, which may be at different integration stages.
@@ 3.1.1. Retriever Type
Retrieval methods can be generally categorized into two types: sparse and dense, based on the information encoding methods. Sparse retrieval is word-based and applied in text retrieval mostly, while dense retrieval embeds queries and external knowledge into vector spaces and can applied to various data formats.
@@
As a straightforward approach, sparse retrieval, e.g., TF-IDF and BM25 (Sparck Jones, 1972; Robertson et al., 2009), usually relies on inverted index matching along with the raw data input. For example, many studies directly apply BM25 for passage-level retrieval to facilitate their RAG (Chen et al., 2017; Ram et al., 2023; Zhong et al., 2022; Jiang et al., 2023; Zhou et al., 2022; Xu et al., 2023b), where passages are specifically represented as a bag of words and ranked based on term and inverse document frequencies (Izacard and Grave, 2021b). On top of offering supplementary to enhance the input of the generator, sparse retrieval has also been used to find demonstrations to function in in-context learning for RA-LLMs (Ye et al., 2023b; Luo et al., 2023a; Rubin et al., 2022; Agrawal et al., 2023; Sia and Duh, 2023). The main limitation of applying sparse retrieval in RAG is its no-training nature, which makes the retrieval performance heavily rely on the quality of the database and the query. Moreover, such fixed term-based methods only support similarity-based retrieval, while cannot be adapted for other retrieval criteria possibly existing in LLM applications, such as the diversity (Drozdov et al., 2022).
@@
Dense retrieval, on the contrary, embeds the query and documents into continuous vector space with certain criteria, for example, semantic similarity (Karpukhin et al., 2020). Dense retrieval methods are usually trainable, therefore holding more flexibility and potential in adaptation. As the key component of dense retriever, the embedding models have delicately different designs in existing RAG models. A simple design (Khandelwal et al., 2020; Lewis et al., 2020a; Wu et al., 2022) is to directly use a part of the generation model as the embedding layer of the retriever, which might be able to enhance the alignment between the retrieval and generation processes. BERT-based backbone (Devlin et al., 2019) is widely applied in retrieval models. One common retriever design in RAG is to construct two-stream encoders with the BERT structure (one encoder for the query and the other for the documents), which is also called bi-encoder (Wu et al., 2020; Shi et al., 2023). Early-stage RAG methods tend to freeze (Borgeaud et al., 2022; Ram et al., 2023) or partially freeze (Lewis et al., 2020c) the parameters of the retriever to perform general-level relevant knowledge extraction and pay more attention to the knowledge leveraging and generator fine-tuning. Large-scale specialized pre-training further enhances RAG models to excel in more knowledge-intensive tasks. One typical success is Dense Passage Retriever (DPR) (Karpukhin et al., 2020), which uses a BERT-based backbone and is pre-trained specifically for the OpenQA task with question-answer pair data. DPR has shown strong capacity as a pre-trained retriever, facilitating many RAG models to succeed in various downstream tasks (Lewis et al., 2020c; Izacard and Grave, 2021b; Siriwardhana et al., 2023; Singh et al., 2021; Shi et al., 2023). It has also been regarded as the first step in the RAG paradigm for improving the performance of LLMs, which may further enhance the alignment of the embeddings between queries and relevant textual data through fine-tuning (Cheng et al., 2023). A recent study (Reichman and Heck, 2024) has also discovered that DPR training decentralizes how knowledge is stored in the network, creating multiple access pathways to the same information. With effective fine-tuning, bi-encoder retrievers are also applied widely in ICL-based RAG (Rubin et al., 2022; Poesia et al., 2022; Lu et al., 2023; Ye et al., 2023b; Milios et al., 2023; Li and Qiu, 2023). Specifically, they have been more often used for sentence embedding similarity-based retrieval, as well as for some special requirement in ICL, such as diverse example retrieval (Ye et al., 2023b).
@@
Another stream of dense retrievers having been widely applied in RA-LLMs uses one encoder only, which may be based on Transformer, BERT or other off-the-shelf sequence modeling backbones. These one-encoder retrievers are generally pre-trained on large-scale unaligned documents by contrastive learning (Reichman and Heck, 2024), which may therefore excel for their versatility, meaning that they can transfer and generalize better to new domains or tasks. Such general-purpose pre-trained retrievers, e.g., Contriever (Gautier et al., 2022) and Spider (Ram et al., 2022), would be more flexible to use in LLMs targeting on various tasks and have demonstrated their effectiveness in many RA-LLM methods, such as In-Context RALM (Ram et al., 2023), Atlas (Izacard et al., 2023) and Self-RAG (Asai et al., 2023b). According to experimental results in existing studies (Yu et al., 2023a), for open-domain QA tasks, when cooperated with InstructGPT (Ouyang et al., 2022), applying general-purpose pre-trained retriever (Contriever) without fine-tuning achieves comparable performance to sparse retriever (BM25). However, they are both worse than the DPR model fine-tuned on target datasets, showing the effectiveness of fine-tuning on targeted tasks and data.